# Gradient Extrapolation for Debiased Representation Learning

### v1

@@ 5.2 Experiments on Datasets-2
We evaluate GERNE on three commonly used datasets[49]: Waterbirds [45], CelebA [30] and CivilComments [5].
@@
We follow the same evaluation strategy in [49] for model selection and hyperparameter tuning: When attributes are known in both training and validation, we use the worst-group accuracy of the test set. When attributes are unknown in training, but known in validation, we use the worst-group accuracy of the validation set. When attributes are unknown in both training and validation, we use the worst-class accuracy of the validation set.
@@
For each dataset, we select the three best performing methods reported in [49]. We end up with ERM[44], Group DRO[39], DFR [18], LISA [50], ReSample [19], Mixup [52], ReWeightCRT [21], ReWeight [19], CBLoss [7], BSoftmax[36] and SqrtReWeight [49]. We also report the results for CnC [53] as it adopts similar training settings.
@@
We employ the same data augmentation techniques, optimizers and pretrained models described in [49]. Further details can be found in Sec. F.2.
@@ Results
Tab. 2 shows the worst-group accuracy of the test set for GERNE against the baseline methods under the evaluation strategy explained above. In the case of known attributes, GERNE achieves the highest performance on the CelebA and CivilComments datasets and ranks second on Waterbirds, following DFR. In case of unknown attributes in training set but known in validation, our approach again achieves the best results on the Waterbirds and CivilComments datasets and remains competitive on CelebA, closely following the top two baseline results. Notably, DFR uses the validation set to train the model, whereas GERNE only uses it for model selection and hyperparameter tuning. We include a comparison of our method’s performance against DFR, where GERNE also uses the validation set for training, in Appendix G. In the scenario where attributes are unknown in both the training and validation sets, our approach achieves the best results on the Waterbirds and CelebA datasets. However, we observe a significant drop in accuracy on CelebA compared to the second case (unknwon attributes in training but known in validation), while this drop is less pronounced on Waterbirds. This can be explained by the worst-class accuracy evaluation metric. In the validation set of CelebA, the majority examples in class 1 exhibit spurious correlations, leading to selection process to favor the majority group while disregarding the minority group. However, the validation set of Waterbirds has balanced groups within each class, resulting in only a slight performance drop for GERNE between the second and third case. This highlights the critical role of having access to the attributes in the validation set or having a group-balanced validation set for model selection and hyperparameter tuning when aiming for better results with GERNE.
@@
Table 2: Performance comparison of GERNE and baseline methods on the Waterbirds, CelebA, and CivilComments datasets. We report the worst-group test accuracy (%) and standard deviation over three trials on the test of each dataset. Baseline results are sourced from [49] as the same experimental settings are adopted. $\checkmark$/$\checkmark$ / ✓ indicates known attributes in training and validation sets. ×$\times$/$\checkmark$ / ✓ in validation set only, and ×/$\times$/$\times$ / × in neither. Best results are highlighted in bold, and the second-best are underlined.
Methods Group Info Waterbirds CelebA CivilComments train attr./val attr. ERM ✓/✓ 69.10 $\pm$ 4.70 62.60 $\pm$ 1.50 63.70 $\pm$ 1.50 Group DRO ✓/✓ 78.60 $\pm$ 1.00 89.00 $\pm$ 0.70 70.60 $\pm$ 1.20 ReWeight ✓/✓ 86.90 $\pm$ 0.70 89.70 $\pm$ 0.20 65.30 $\pm$ 2.50 ReSample ✓/✓ 77.70 $\pm$ 1.20 87.40 $\pm$ 0.80 73.30 $\pm$ 0.50 CBLoss ✓/✓ 86.20 $\pm$ 0.30 89.40 $\pm$ 0.70 73.30 $\pm$ 0.20 DFR ✓/✓ 91.00 $\pm$ 0.30 90.40 $\pm$ 0.10 69.60 $\pm$ 0.20 LISA ✓/✓ 88.70 $\pm$ 0.60 86.50 $\pm$ 1.20 73.70 $\pm$ 0.30 GERNE (ours) ✓/✓ 90.20 $\pm$ 0.22 91.98 $\pm$ 0.15 74.65 $\pm$ 0.20 ERM $\times$/✓ 69.10 $\pm$ 4.70 57.60 $\pm$ 0.80 63.20 $\pm$ 1.20 Group DRO $\times$/✓ 73.10 $\pm$ 0.40 78.50 $\pm$ 1.10 69.50 $\pm$ 0.70 ReWeight $\times$/✓ 72.50 $\pm$ 0.30 81.50 $\pm$ 0.90 69.90 $\pm$ 0.60 DFR $\times$/✓ 89.00 $\pm$ 0.20 86.30 $\pm$ 0.30 63.90 $\pm$ 0.30 Mixup $\times$/✓ 78.20 $\pm$ 0.40 57.80 $\pm$ 0.80 66.10 $\pm$ 1.30 LISA $\times$/✓ 78.20 $\pm$ 0.40 57.80 $\pm$ 0.80 66.10 $\pm$ 1.30 BSoftmax $\times$/✓ 74.10 $\pm$ 0.90 83.30 $\pm$ 0.30 69.40 $\pm$ 1.20 ReSample $\times$/✓ 70.00 $\pm$ 1.00 82.20 $\pm$ 1.20 68.20 $\pm$ 0.70 CnC $\times$/✓ 88.50$\pm$ 0.30 88.80$\pm$ 0.90 68.90$\pm$ 2.10 GERNE (ours) $\times$/✓ 90.21 $\pm$ 0.42 86.28 $\pm$ 0.12 71.00 $\pm$ 0.33 ERM $\times$/$\times$ 69.10 $\pm$ 4.70 57.60 $\pm$ 0.80 63.20 $\pm$ 1.20 Group DRO $\times$/$\times$ 73.10 $\pm$ 0.40 68.30 $\pm$ 0.90 61.50 $\pm$ 1.80 DFR $\times$/$\times$ 89.00 $\pm$ 0.20 73.70 $\pm$ 0.80 64.40 $\pm$ 0.10 Mixup $\times$/$\times$ 77.50 $\pm$ 0.70 57.80 $\pm$ 0.80 65.80 $\pm$ 1.50 LISA $\times$/$\times$ 77.50 $\pm$ 0.70 57.80 $\pm$ 0.80 65.80 $\pm$ 1.50 ReSample $\times$/$\times$ 70.00 $\pm$ 1.00 74.10 $\pm$ 2.20 61.00 $\pm$ 0.60 ReWeightCRT $\times$/$\times$ 76.30 $\pm$ 0.20 70.70 $\pm$ 0.60 64.70 $\pm$ 0.20 SqrtReWeight $\times$/$\times$ 71.00 $\pm$ 1.40 66.90 $\pm$ 2.20 68.60 $\pm$ 1.10 CRT $\times$/$\times$ 76.30 $\pm$ 0.80 69.60 $\pm$ 0.70 67.80 $\pm$ 0.30 GERNE (ours) $\times$/$\times$ 89.88 $\pm$ 0.67 74.24 $\pm$ 2.51 63.10 $\pm$ 0.22
@@
Balancing techniques have been shown to achieve state-of-the-art results, while remaining easy to implement [16, 49]. Notably, resampling methods often outperform reweighting strategies when combined with stochastic gradient algorithms [2]. Our results, as presented in Tab. 1, demonstrate that GERNE outperforms resampling (GERNE with $c=1,\beta=0$) when the evaluation metric is Group-Balanced Accuracy or Accuracy on minority group. More discussion in Appendix B. This highlights the flexibility of GERNE to adapt to maximize both metrics, and its superior performance in comparison to resampling and other balancing techniques, as further supported by the results in Tab. 2. In Appendix C, we provide a detailed ablation study that compares GERNE to an equivalent (in term of loss expectation)’sampling + weighting’ approach, and show how GERNE can leverage its controllable loss variance (tuned by the hyperparameters $c,\beta$) to escape sharp minima.
@@ 5.4 Ablation Study
The value of $\beta$ in Eq. 7 has a significant impact on the performance of our method in debiasing the model (i.e., leading the training process in a debiased direction and avoid learning spurious features). In Fig. 2, we show how tuning $\beta$ affects the learning process in the case of the C-MNIST dataset with $0.5\%$ of minority group in the known attribute case. We show results for $\beta\in\{-1,0,1,1.2\}\text{ with }c=0.5$. For $\beta=-1$, our target loss $\mathcal{L}_{ext}$ in Eq. 5 equals the biased loss $\mathcal{L}_{b}$, which leads to learning a biased model that exhibits high accuracy in the majority group, yet demonstrates poor performance on both the minority group and the unbiased test set. As $\beta$ increases (e.g. $\beta=0,\beta=1$), the model starts learning more intrinsic features. This is evident from the improved performance on the minority group in the validation set, as well as on the unbiased test set. However, as the extrapolation factor $\beta$ continues to increase, the model begins to exhibit higher variance during the training process as shown for $\beta=1.2$, ultimately leading to divergence when $\beta$ exceeds the upper bound defined in Eq. 9 which is equal to $1.22$ in this case. By comparing the accuracies of Minority/Majority training groups in case $\beta=0,\beta=1$, we can see that both cases have around $100\%$ accuracy on minority but higher accuarcy on majority for $\beta=0$. However, we notice a better generalization when $\beta=1$. This highlights the importance of directing the training process to the right direction early in training while overfitting is expected as well.
Figure 2: The impact of tuning $\beta$ in debiasing the model $\beta\in\{-1,0,1,1.2\}$ on debiasing the model. On the left column, we plot the training losses $\mathcal{L}_{b},\mathcal{L}_{lb}$ and the target loss $\mathcal{L}_{ext}$. On the right column, we plot the average accuracy of the minority and majority groups in both training and validation, as well as the average accuracy of the unbiased test set. Each plot represents the mean and standard deviation calculated over three runs with different random seeds.
@@
To answer this question, we conduct experiments on C-MNIST dataset with $0.5\%$ of minority group. We first train a biased model $\tilde{f}$, and use its predictions to generate the pseudo-attributes for five different values of the threshold $t$. Let’s refer to the pseudo-groups with $\tilde{a}=1$ as the pseudo-minority group. For each threshold, we tune $\beta$ to achieve the best average accuracy on test set. Simultaneously, we compute the average precision and recall for the minority group, as defined in Eq. 14. As shown in Fig. 3, with $t=0.0005$, the average precision reaches 1, indicating that all the samples in the pseudo-minority group are from the minority group. However, these samples constitute less than 20% of the total number of samples in the minority group, as indicated by the average recall. Despite this, GERNE achieves a high accuracy of approximately 70%, remaining competitive with other methods reported in Tab. 1 while using only a very limited number of minority samples ($t=$0.0005 corresponds to about 28 samples versus 249 minority samples out of 55,000 samples in the training set). As $t$ increases to 0.001 and 0.003, precision remains close to 1 while increasing the number of minority samples in the pseudo-minority group. This increase introduces more diversity among minority samples within the pseudo-minority group, allowing for lower $\beta$ values to achieve the best average accuracy on test set. However, for even higher thresholds, such as $t=$0.01, minority samples constitute less than 40% in the pseudo-minority group, prompting a need to revert to higher $\beta$ values. We conclude that identifying the minority group is of utmost importance for achieving optimal results (high average precision and high recall) and this agrees with the results presented in both Tab. 1, Tab. 2 where we achieve the best results in the case of known attributes.
Figure 3: The effect of choosing the threshold $t$ to generate the pseudo attributes on $\beta$ for best performance.
Average Precision $=\frac{1}{K}\sum_{y\in\mathcal{Y}}\frac{\#\text{ minority in }(y,\tilde{a}=1)}{\#\text{ total in }(y,\tilde{a}=1)},$ (14)
Average Recall $=\frac{1}{K}\sum_{y\in\mathcal{Y}}\frac{\#\text{ minority in }(y,\tilde{a}=1)}{\#\text{ minority in }(y)}$
@@ 6 Conclusion
In this work, we introduce GERNE, a novel debiasing approach designed to mitigate spurious correlations. By sampling two types of batches with varying amounts of spurious correlations, we define a debiasing direction using the gradients of the losses computed on these two batches. The extrapolated gradient vector, controlled by the extrapolation factor, is then used to update the model parameters. We demonstrate that GERNE can be adapted to optimize the Group-Balanced Accuracy (GBA) metric or the Worst-Group Accuracy (WGA) metric, leading to improved generalization and superior performance compared to state-of-the-art methods on both known and unknown attribute cases, without relying on data augmentation techniques, on multiple benchmarks. Notably, GERNE provides a general framework that encompasses common methods such as ERM and resampling, making it a valid approach for training models even on unbiased datasets. Furthermore, GERNE allows for control over the degree of loss variance during training, increasing the likelihood of escaping sharp minima and consequently improving generalization. Although effective, GERNE remains limited in its utilization of the diversity within the majority group, where spurious correlations are present. This raises a new research question: Can we identify a more effective debiasing direction based on the two previously mentioned gradients? Future work will aim to address this question. Additionally, we will explore strategies to dynamically update the extrapolation factor during training to reduce reliance on hyperparameter tuning, and investigate novel methods to better estimate the attributes in the unknown attribute case.
@@
Appendix A Conditional attribute distribution of the extrapolated loss: Proof of Eq. 8
@@
[…]
Where $p_{b}(a|y),p_{lb}(a|y)$ defined in Eq. 2, Eq. 4, respectively. Substituting Eq. 15, Eq. 16 in Eq. 5:
@@
Appendix B GERNE versus the resampling method [16]
In Tab. 1, GERNE achieves higher GBA compared to the special case of GERNE with $c=1,\beta=0$ (resampling method) for both C-MNIST and C-CIFAR-10 datasets. Our explanation behind GERNE superior performance is that resampling method tend to present the majority and minority groups equally in the batch, and the model $f$ tends to prioritize learning the easy-to-learn spurious features associated with the majority group. For instance, the color in C-MNIST. While GERNE undermines learning the spurious correlations by directing the learning process more in the debiasing direction thanks to the extrapolation factor.
Appendix C GERNE versus an equivalent sampling and weighting approach
We compare GERNE with an equivalent (in term of loss expectation) sampling+weighting method, which we refer to as ’SW’. For simplicity, we assume the following:

### v2

@@ 5.2 Experiments on Datasets-2
Waterbirds  [46] contains bird images with spurious correlations between bird type and background: Most waterbirds appear with water backgrounds, while most landbirds appear with land backgrounds. CelebA  [30] involves classifying hair color (blond, non-blond), with gender (male, female) as the spurious attribute: Most blond images depict females. CivilComments  [5] is a binary toxic comment classification dataset, where the spurious attribute marks references to eight different demographic identities (male, female, LGBTQ, Christian, Muslim, other religions, Black, and White).
@@
We follow the same evaluation strategy from  [50] for model selection and hyperparameter tuning. When attributes are known in both training and validation, we use the worst-group test accuracy as the evaluation metric. When attributes are unknown in training, but known in validation, we use the worst-group validation accuracy. When attributes are unavailable in both, we use the worst-class validation accuracy.
@@
For each dataset, we select the best three methods reported in  [50]. We end up with ERM [45], Group DRO [39], DFR  [18], LISA  [51], ReSample  [19], Mixup  [53], ReWeightCRT  [21], ReWeight  [19], CBLoss  [7], BSoftmax [36] and SqrtReWeight  [50]. We also report the results for CnC  [54] as it adopts similar training settings.
@@
We employ the same data augmentation techniques, optimizers and pretrained models described in  [50]. Further details are in Sec. F.2.
@@ Results
Tab. 2 shows the worst-group accuracy (WGA) of the test set for GERNE compared to the baseline methods under the evaluation strategy explained above. In the known attributes case, GERNE achieves the highest accuracy on CelebA and CivilComments, and ranks second on Waterbirds, following DFR. In case of unknown attributes in the training set but known in validation, our approach again attains the best results on Waterbirds and CivilComments datasets and remains competitive on CelebA, closely following the top two baselines’ results. In particular, DFR uses the validation set to train the model, whereas GERNE employs it only for model selection and hyperparameter tuning. We include a comparison between DFR and GERNE when using the validation set for training in Appendix G. When attributes are unknown in both the training and validation sets, GERNE achieves the best results on Waterbirds and CelebA. However, we observe a significant drop in accuracy on CelebA compared to the second case (known attributes only in validation), while this drop is less pronounced on Waterbirds. The difference can be attributed to the use of worst-class accuracy as the evaluation metric. In CelebA’s validation set, the majority of blond hair images exhibit spurious correlations (female images), leading the model selection process to favor the majority group while disregarding the minority group. In contrast, the validation set in Waterbirds is group-balanced within each class, leading to only a slight decrease in performance between the second and third case. This highlights the critical role of having access to the attributes in the validation set —or at least a group-balanced validation set— for model selection when using GERNE.
@@
Table 2: Performance comparison of GERNE and baseline methods on Waterbirds, CelebA, and CivilComments. We report the worst-group test accuracy (%) and standard deviation over three trials for each dataset. Baseline results are sourced from  [50] as the same experimental settings are adopted. $\checkmark/\checkmark$ denotes known attributes in training and validation sets. $\times/\checkmark$ indicates attributes are known only in the validation set, while $\times/\times$ signifies that attributes are unknown in both sets. Best results are highlighted in bold, and the second-best are underlined.
[…]
@@
Balancing techniques have been shown to achieve state-of-the-art results, while remaining easy to implement  [16, 50]. While Resampling often outperforms Reweighting when combined with stochastic gradient algorithms [2], we show in Tab. 1 that GERNE consistently outperforms Resampling in both group-balanced accuracy (GBA) and minority group accuracy. This highlights the flexibility of GERNE to adapt to maximize both metrics, and its superior performance in comparison to Resampling and other balancing techniques, as further supported by the results in Tab. 2. In Appendix B, we provide a detailed ablation study comparing GERNE to an equivalent “sampling+weighting” approach with matching loss expectation, and demonstrate how GERNE can leverage its controllable loss variance (by the hyperparameters $c,\beta$) to escape sharp minima.
@@ 5.4 Ablation Study
The value of $\beta$ in Eq. 7 plays a critical role in guiding the model toward learning debiased representation (i.e., reducing reliance on spurious features and improving generalization). In Fig. 2, we illustrate the effect of tuning $\beta$ on the learning process using C-MNIST with $0.5\%$ of minority group in the known attributes case. We show results for $\beta\in\{-1,0,1,1.2\}\text{ with }c=0.5$. For $\beta=-1$, our target loss $\mathcal{L}_{ext}$ in Eq. 5 equals the biased loss $\mathcal{L}_{b}$, which leads to learning a biased model that exhibits high accuracy on the majority group, yet demonstrates poor performance on both the minority group and the unbiased test set. As $\beta$ increases (e.g., $\beta=0,\beta=1$), the model starts learning more intrinsic features. This is evident from the improved performance on the minority group in the validation set, as well as on the unbiased test set. However, as the extrapolation factor $\beta$ continues to increase, the model begins to exhibit higher variance during the training process, as shown for $\beta=1.2$, ultimately leading to divergence when $\beta$ exceeds the upper bound defined in Eq. 9 ($1.22$ in this case). While GERNE appears to be sensitive to small variations in $\beta$ (e.g., 1.2 to 1.22), we show in Appendix C that $\beta_{\max}$ is inversely proportional to $c$ and $A$. This implies that decreasing $c$ allows for a wider feasible range of $\beta$. By comparing the accuracies on minority and majority training groups in case $\beta=0,\beta=1$, we can see that both cases have around $100\%$ accuracy on minority but higher accuracy on majority for $\beta=0$. However, $\beta=1$ results in better generalization overall. This highlights the importance of directing the training process toward a debiased direction early in training, especially when overfitting is likely to occur on the minority group (e.g., when it contains very few samples).
Figure 2: The impact of tuning $\beta\in\{-1,0,1,1.2\}$ on debiasing the model. On the left column, we plot the training losses $\mathcal{L}_{b},\mathcal{L}_{lb}$ and the target loss $\mathcal{L}_{ext}$. On the right column, we plot the average accuracy of the minority and majority groups in both training and validation sets, as well as the average accuracy of the unbiased test set. Each plot represents the mean and standard deviation calculated over three runs with different random seeds.
@@
To answer this question, we conduct experiments on C-MNIST dataset with $0.5\%$ of minority group. We first train a biased model $\tilde{f}$, and use its predictions to generate the pseudo-attributes for five different values of the threshold $t$. Let’s refer to the pseudo-groups with $\tilde{a}=1$ as the pseudo-minority groups. For each threshold, we tune $\beta$ to achieve the best average test accuracy. Simultaneously, we compute the average precision and recall for the minority group. As shown in Fig. 3, with $t=5\times 10^{-4}$, the average precision reaches 1, indicating that all the samples in the pseudo-minority groups are from the minority groups. However, these samples constitute less than 20% of the total minority groups, as indicated by the average recall. Despite this, GERNE achieves a high accuracy of approximately 70%, remaining competitive with other methods reported in Tab. 1 while using only a very limited number of minority samples ($t=5\times 10^{-4}$ corresponds to about 28 samples versus 249 minority samples out of 55,000 samples in the training set). As $t$ increases to $10^{-3}$ and $3\times 10^{-3}$, precision remains close to 1 while increasing the number of minority samples in the pseudo-minority groups. This increase introduces more diversity among minority samples within the pseudo-minority groups, allowing for lower $\beta$ values to achieve the best average test accuracy. However, for even higher thresholds, such as $t=10^{-2}$, minority samples constitute less than 40% in the pseudo-minority groups, prompting a need to revert to higher $\beta$ values. We conclude that identifying the samples of minority groups (high average precision and high recall) is of utmost importance for achieving optimal results, and this agrees with the results presented in both Tab. 1, Tab. 2, where we achieve the best results in the known attributes case.
Figure 3: The effect of the threshold $t$ used to generate pseudo-attributes on the extrapolation factor $\beta$ and model performance. We plot the average precision and recall over pseudo-minority groups $(y,\tilde{a}=1)$, averaged across all classes $y$. For each $(y,\tilde{a}=1)$, precision is defined as the fraction of minority samples among all samples in that group, and recall is the fraction of those minority samples relative to all minority samples in class $y$. We also report the best achievable test accuracy, along with the corresponding extrapolation factor $\beta$, across different threshold values.
@@ 6 Conclusion
We introduce GERNE, a novel debiasing approach that effectively mitigates spurious correlations by leveraging an extrapolated gradient update. By defining a debiasing direction from loss gradients computed on batches with varying degrees of spurious correlations, GERNE’s tunable extrapolation factor allows optimizing either Group-Balanced Accuracy (GBA) or Worst-Group Accuracy (WGA). Our comprehensive evaluations across vision and NLP benchmarks demonstrate GERNE’s superior performance over state-of-the-art methods, both for known and unknown attribute cases, without data augmentation. Furthermore, GERNE offers a general framework that encompasses methods like ERM and Resampling, extending its applicability to unbiased datasets. Future work will explore dynamic adaptation of the extrapolation factor and refine attribute estimation for the unknown attributes case.
Acknowledgments.
This work was funded by the Carl Zeiss Foundation within the project Sensorized Surgery, Germany (P2022-06-004). Maha Shadaydeh is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Individual Research Grant SH 1682/1-1.
@@ Acknowledgments.
Appendix A The Extrapolated Conditional Attribute Distribution of GERNE:
@@
$\mathcal{L}_{b}=\int\ell(y,f(x))p(y)\,p_{b}(a|y)\,p(x|a,y)\,dx\,dy\,da,$ (14)
$\mathcal{L}_{lb}=\int\ell(y,f(x))p(y)\,p_{lb}(a|y)\,p(x|a,y)\,dx\,dy\,da.$ (15)
Where $p_{b}(a|y),p_{lb}(a|y)$ defined in Eq. 2, Eq. 4, respectively, and $x$ is uniformly sampled within each group $\mathcal{X}_{y,a}$ (i.e., $p(x|y,a)=\frac{1}{|\mathcal{X}_{y,a}|}$). Substituting Eq. 14, Eq. 15 in Eq. 5:
@@
Furthermore, we can write $\mathcal{L}_{ext}$ as follows:
$\mathcal{L}_{\text{ext}}$ $=\mathbb{E}_{y\sim p(y)}\left[\mathbb{E}_{a\sim p_{\text{ext}}(a|y)}\left[\mathbb{E}_{x\sim p(x|a,y)}\left[\ell(y,f(x))\right]\right]\right]$
$=\frac{1}{K}\cdot\sum_{g=(y,a)\in\mathcal{G}}p_{\text{ext}}(a|y;\beta)\cdot L_{g},$
with $L_{g}=\mathbb{E}_{x\sim p(x|(y,a)=g)}\left[\ell(y,f(x))\right]$ by using the discrete expectations over $y$ and $a|y$, with $p(y)=\frac{1}{K}$.
Appendix B GERNE Versus an Equivalent Sampling and Weighting Approach
We compare GERNE with an equivalent (in terms of loss expectation) sampling+weighting method, which we refer to as “SW”. For simplicity, we assume the following: